# Simulations of gravitational collapse in null coordinates IV: evolving through the event horizon, with an application to the spherical charged scalar field

### v1

@@ Appendix C Expansions at the centre
$\ln\left({G\over R_{,x}}\right)$ $=$ $2\pi(\psi_{(1)}^{2}+\chi_{(1)}^{2})x^{2}$ (118)
@@ Appendix D Radial null geodesics peeling off the horizon
We have so far defined $\bar{v}$ only on RN, not on the collapse spacetime that we are trying to model. We note that on RN $\bar{v}$ obeys $\bar{v}=\bar{u}+2R$. If we assume that for sufficiently early $u$ and sufficiently late $\bar{v}$ the collapse spacetime is approximated by the RN exterior, we can characterize $\bar{v}$ on the collapse spacetime as the retarded time that obeys $\left.R_{,\bar{v}}\right|_{u}=2$ in that region. We could then define $\bar{v}=2R$ on $u=0$ and propagate it with $\Xi\bar{v}=0$.
@@ Appendix E Boundary conditions on an ingoing null cone intersecting an outgoing one
where $C_{l}$ is used as a shorthand for
@@
We now have the data that we need to start up the integration of the hierarchy equations at $x=0$, namely $Q=Q_{l}$ for (27), $A=0$ for (30), either $R=R_{l}$ and $V=V_{l}$ for (64,65) or $G=G_{*}$ for (39), $\Xi R=\dot{R}_{l}$ for (42), $\hat{\Xi}\psi=\dot{\psi}_{l}$ for (43), $\hat{\Xi}\chi=\dot{\chi}_{l}$ for (44), and ${\cal H}=0$ for (45).
@@ Appendix F Discretization of the left boundary equations
The Runge-Kutta ODE solver that we use for the time integration assumes that the ODE system is autonomous, we add $u$ as an evolved variable to the set of left-boundary variables, in order to access the given functions $\dot{\psi}_{l}(u)$ and $\dot{\chi}_{l}(u)$ at the correct substep values of $u$.
@@ Appendix G Convergence testing
$f(u,x;h)=f_{0}(u,x)+h^{2}f_{2}(u,x)+h^{3}f_{3}(u,x)+O(h^{4}),$ (141)
where $f_{0}(u,x)$ is the (unknown) continuum solution and $f_{i}(u,x)$ are (discretisation-dependent, unknown) error terms, and hence

### v2

@@ Appendix C Expansions at the centre
$\ln\left({G\over R_{,x}}\right)$ $=$ $2\pi(\psi_{(1)}^{2}+\chi_{(1)}^{2})x^{2}$
@@ Appendix D Radial null geodesics peeling off the horizon
We have so far defined $\bar{v}$ only on RN, not on the collapse spacetime that we are trying to model. We note that on RN $\bar{v}$ obeys $\bar{v}=\bar{u}+2R$. If we assume that for sufficiently early $u$ and sufficiently late $\bar{v}$ the collapse spacetime is approximated by the RN exterior, we can define $\bar{v}=2R$ on $u=0$ and propagate it with $\Xi\bar{v}=0$.
@@ Appendix E Boundary conditions on an ingoing null cone intersecting an outgoing one
where $R_{,xl}(u):=R_{,x}(u,0)$ and $C_{l}$ is used as a shorthand for
@@
We now have the data that we need to start up the integration of the hierarchy equations at $x=0$, namely $Q=Q_{l}$ for (27), $A=0$ for (30), either $R=R_{l}$ and $V=V_{l}$ for (65,64) or $G=G_{*}$ for (39), $\Xi R=\dot{R}_{l}$ for (42), $\hat{\Xi}\psi=\dot{\psi}_{l}$ for (43), $\hat{\Xi}\chi=\dot{\chi}_{l}$ for (44), and ${\cal H}=0$ for (45).
@@ Appendix F Discretization of the left boundary equations
The Runge-Kutta ODE solver that we use for the time integration assumes that the ODE system is autonomous. We add $u$ as an evolved variable to the set of left-boundary variables, in order to access the given functions $\dot{\psi}_{l}(u)$ and $\dot{\chi}_{l}(u)$ at the correct substep values of $u$.
@@ Appendix G Convergence testing
$f(u,x;h)=f_{0}(u,x)+h^{2}f_{2}(u,x)+h^{3}f_{3}(u,x)+O(h^{4})$ (141)
to hold, where $f_{0}(u,x)$ is the (unknown) continuum solution and $f_{i}(u,x)$ are (discretisation-dependent, unknown) error terms, and hence